# Learning Successor Features with Distributed Hebbian Temporal Memory

### v3

@@ 3.3 Learning
DHTM learns $f_{l}$ and $w_{ul}$ weights by Monte-Carlo Hebbian-like updates. First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: ${p(h^{i}_{t-1}\ |\ \varphi^{i}_{t-1})\propto M^{i}_{t-1}}$ and ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ respectively. Then $f_{l}$ is updated according to the activity $s_{l}$ of the segment and the activity $c^{j}_{t}$ of its cell, so that $f_{l}$ is proportional to several coincidences ${s_{l}=c^{j}_{t}=1}$ in the recent past:
@@
That is, $f_{l}$ is trained to be equal to the average cell activation frequency when the segment is active. It’s similar to Baum-Welch’s update rule [41] for the transition matrix in HMM, which in effect counts transitions from one state to another, but in our case, the previous state (context) is represented by a group of RVs.
Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$. That is, they are targeted to represent the probability ${p(s_{l}=1\ |\ c^{u}_{t-1}=1)}$ that segment $s_{l}$ is active, given that cell $u$ was active at the previous time step. We could learn this by counting activation matches and mismatches. But in our algorithm, it is approximated as an exponential moving average of the frequency activation of segment $s_{l}$, given $c^{u}_{t-1}=1$:
@@ 3.4 Agent Architecture
4:     while (not terminal) (steps $<$ max_steps) do
@@
The agent training procedure is outlined in Algorithm 1. For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and the variable action is set to a fixed initial action. An observation returned by an environment (obs) is encoded as a set of categorical variables. In the OBSERVE() routine, memory learns to predict next feature states as described in section 3. An agent learns associations to feature states and rewards in the REINFORCE() function:
@@
In SAMPLE_ACION(), for each action and current state, TM predicts feature states up to time step $T$ under policy $\pi$, and the predictions are combined to form SF (equation 11). For simplicity, we always build SF under a uniform policy, which doesn’t guarantee optimality, but we found that it doesn’t make a significant difference to iterative policy improvement in our experimental setups. The planning horizon $T$ is determined dynamically from the distribution of $\Phi^{k}_{t+l}$. If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with positive reward have high probability, the prediction cycle is stopped. The action value is determined by combining the corresponding SF and feature state rewards as shown in equation 12. Finally, the action is sampled from the softmax distribution over the action values.
@@ 3.5 DHTM and Episodic Control
DTHM performs best when its hidden state space is large. Since the hidden state for unpredictable observations is formed randomly, it’s important that there are no accidental state collisions when remembering transitions. That is, DHTM tries to remember new sequences without interfering with old memories. In effect, this makes DHTM a buffer of agent trajectories with a built-in recall engine. Episodic Control is based on a similar idea, and to investigate its difference from DHTM, we implemented an Episodic Control agent, a simplified counterpart of DHTM, reflecting its core property that presumably helps it solve RL tasks.
In our EC version, $K_{a}$ is an array of hidden states $h_{t}$ and $V_{a}$ is an array of corresponding subsequent hidden states $h_{t+1}$. If the key state $h_{t}$ already exists in the dictionary, its value is rewritten. That is, $\mathcal{D}_{a}$ represents the transition matrix of a deterministic HMM for a feature variable $\Phi_{t}$. The hidden state $h_{t}$ and the feature state $\varphi_{t}$ have the same dependence as in DHTM (equation 5), but here we assume infinite cells per column. To predict the next state, we just have to look up $\mathcal{D}_{a}$ for the current state $h_{t}$, but if there is no entry for $h_{t}$ or the prediction does not match the observed state $\varphi_{t+1}$, then the new state is formed. Unlike DHTM, the new state is not completely random, but to avoid collisions, the algorithm makes sure that this state hasn’t been chosen before. The learning procedure of the EC agent’s memory $\mathcal{D}_{a}$ is summarized in the algorithm 2.
Since $\mathcal{D}_{a}$ can be transformed into a transition matrix, the process of forming SF can be reduced to one in Sec. 3.4. However, since the resulting transition matrix is sparse, it’s more efficient to consider $\mathcal{D}_{a}$ as a graph representation and to perform SF formation as a breadth-first search (see Appendix D.3 for details).
@@ 4 Experiments
We test our model in a reinforcement learning task in the Gridworld environment and in a more challenging AnimalAI 3D environment. This section shows how different memory models affect the adaptability of an RL agent. For each setup, the results are averaged over five experiments with different seed values, and the standard deviation is shown. See Appendix D for details on the baselines, their model parameters, and training regimes. See Appendix E for a more detailed description of the environments and setups. The source code of the experiments and algorithms is at https://anonymous.4open.science/r/dhtm-FA0E.
@@ 4.1 Gridworld
Figure 2: POMDP 5x5 Gridworld foraging task using different types of memory: DHTM, CSCG, LSTM. CSCG and LSTM are trained in two modes. The first mode is with oracle, which tells the agent to reinitialize memory and clear the trajectory buffer during environmental change. The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories and update memory. There is also an episodic control agent (ec) among baselines, which uses a simple dictionary to store trajectories and make predictions. A. Dynamic of steps required for the agent to reach the goal. In the 300th episode, the goal is blocked by a wall, so the agent needs to find a new optimal trajectory. B. Total steps taken during 1000 episodes.
@@
The first test was conducted in partially observable Gridworld in a 5x5 alternating maze with a stationary end state and the agent’s starting position. The agent can observe the colour of the cell it is in and the colour of an obstacle it encounters, while being punished in addition to the constant step penalty. Since the observation can be represented directly as a categorical variable with the number of states equal to the number of possible colours, no encoder is needed. Each episode starts with the agent in the same position and ends when the agent enters the terminal state (i.e., when the reward is collected) or when the maximum number of action steps is reached. In each episode, we measure the number of action steps required for the agent to reach the goal.
@@
Each trial lasts $1000$ episodes and consists of two phases. In the first phase, which lasts $300$ episodes, there are no obstacles on the shortest path to the terminal state. In the second phase, the previously optimal path is blocked by a wall, so the agent has to take a detour to get the reward. The experiment aims to show the ability of a memory to relearn the structure of the maze in the test time. If the agent is able to reach the end state in the second phase, it suggests that the model can forget, or at least suppress, the memory of the old shortest path when it is blocked. Otherwise, the agent will continue to hit the wall and fail to reach the terminal state.
@@
In this experiment, we compare our memory model DHTM with CSCG [39] and LSTM [9] within the same agent architecture that uses memory to form successor features and evaluate actions (see Sec. 3.4). CSCG and LSTM are trained in two modes. The first mode is with an oracle that tells the agent to reinitialize memory and clear the trajectory buffer just in time for the environmental change. The second mode is without an oracle, so the agent continues to use its current strategy to collect new trajectories and update memory in the second phase. There is also a simple episodic control agent (EC) among the baselines, which uses a dictionary to store trajectories and make predictions (see Section 3.5 for details).
@@
The results in Figure 2 show that CSCG effectively can’t adapt to the environmental change in the online regime, but only if it is retrained from scratch on new trajectories. And, apparently, LSTM is only able to learn a predictive model for a small horizon, since it performs significantly worse than other baselines in the second phase, even with oracle enabled. In contrast, DHTM and EC are able to adapt online without explicit memory reset. Another important observation is that the EC agent performs on par with DHTM, suggesting that its mechanism can explain most of DHTM’s performance in this task.
@@ 4.2 AnimalAI
Figure 3: Amount of action steps the agent performs to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM, CSCG, LSTM and EC dictionary. The agent and the goal have the same initial positions throughout the episodes. Each time-step, the agent gets an encoded image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward. A. Dynamic of steps required for the agent to reach the goal. B. Total steps taken during 1000 episodes by the agent with DHTM with K-Means encoder and Categorical VAE. For the latter, DHTM’s context factor size is varied (i.e., how many previous time-step hidden variables are considered for prediction).
@@
We test the agent using two different encoders: K-Means and Categorical VAE [46]. Both encoders are pre-trained on RGB images collected with a uniform strategy during $2000$ episodes in the same room used for agent testing. K-Means represents an image as a single categorical variable with $50$ states, and Categorical VAE is configured for distributed representation with five variables with $50$ states each (see Appendix C for details).
@@
To show that DHTM can’t simply be replaced by multiple parallel EC dictionaries or CSCGs in the case of distributed encoding, we also vary the context factor size of DHTM, that is, how many previous time-step variables are taken into account to predict the current state of a variable. The factor size of 1 corresponds to the graph of the Factorial HMM, with parallel HMMs. It can be seen from Figure 3B that in our case of categorical VAE, apparently knowing only the previous state of the variable is not enough to reliably predict rewarding features, since the agent doesn’t converge to the shortest path. We also tested the variant (factor 1*) where each feature variable is represented by three hidden states, to ensure that the result for factor size $1$ is not due to DHTM’s hidden state collisions. DHTM with a factor that takes into account more than one previous time-step hidden variable (excluding the action variable) performs significantly better. For diagrams depicting factor graphs used in these experiments, see Appendix D.4.
@@ 5 Conclusion
Our results suggest that DHTM can be viewed as an adaptive trajectory buffer for a slower generalizing algorithm, such as a Neural State Space Model-based [47]. Through prediction error learning, DHTM can selectively discard and replace obsolete transitions in observed trajectories without disrupting relevant experience. It can also provide an initial policy for an agent through episodic control until a slower but more accurate world model takes over. These features of a trajectory buffer are highly desirable in the case of online learning in non-stationary environments.
@@ Appendix A Value Function Decomposition
Appendix B Biological Interpretation
Neural implementation of the DHTM is inspired by neocortical neural networks (see Fig. 4). Hidden variables $H^{k}$ may be considered as populations of excitatory pyramidal neurons in cortical layer L2/3 of somatosensory areas, with lateral inhibition modelled as $\mathrm{softmax}$ function. Staiger and Petersen [48] showed that neurons in this layer are responsible for temporal context formation.
The neuronal activity at timestep $t$ can be thought to carry messages $M^{k}_{t-1}$. Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$. Dendritic segments of biological neurons are known to be coincidence detectors of their synaptic input [45]. We use the notion of dendritic segment to sparsely represent context factors $F^{k}_{c}$, as each factor value corresponds to a particular combination of states (or active cells).
Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3. L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [49].

### v4

@@ 3.3 Learning
DHTM learns $f_{l}$ and $w_{ul}$ weights by Monte-Carlo Hebbian-like updates. First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: ${p(h^{i}_{t-1}\ |\ \varphi^{i}_{t-1})\propto M^{i}_{t-1}}$ and ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ respectively. Cells that correspond to sampled states become active. For cell $j$ corresponding to $h^{k}_{t}$ we check if there are segments matching the pattern of sampled states (active cells) from the previous time-step $\{h^{i}_{t-1}\}_{i}$. If there are no such segments, the new segment connecting sampled cells is created. For all newly created segments and segments with receptive fields matching active cells from the previous time-step, its activity $s_{l}$ is set to $1$. Then $f_{l}$ is updated according to the activity of the segment $s_{l}$ and the activity of its cell $c^{j}_{t}$, so that $f_{l}$ is proportional to several coincidences ${s_{l}=c^{j}_{t}=1}$ in the recent past (exponential moving average of coincidence):
@@
That is, $f_{l}$ is trained to be equal to the cell’s average activation frequency when the segment is active. It’s similar to Baum-Welch’s update rule (Baum et al., 1970) for the transition matrix in HMM, which, in effect, counts transitions from one state to another, but in our case, the previous state (context) is represented by a group of RVs.
Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$. That is, they are targeted to represent the probability ${p(s_{l}=1\ |\ c^{u}_{t-1}=1)}$ that segment $s_{l}$ is active, given that cell $u$ was active at the previous time step. In our algorithm, it is approximated as an exponential moving average of the frequency activation of segment $s_{l}$, given $c^{u}_{t-1}=1$:
@@ 3.4 Agent Architecture
4:     while (not terminal) and (steps $<$ max_steps) do
@@
The agent training procedure is outlined in Algorithm 1. For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and the variable action is set to a fixed initial action. An observation returned by an environment (obs) is encoded as a set of categorical variables. In the OBSERVE() routine, memory learns to predict next feature states as described in section 3. An agent learns associations to map feature states and rewards in the REINFORCE() function:
@@
In SAMPLE_ACION(), for each action and current state, TM predicts feature states up to time step $T$ under policy $\pi$, and the predictions are combined to form SF (equation 10). For simplicity, we always build SF under a uniform policy, which doesn’t guarantee optimality, but we found that it doesn’t make a significant difference to iterative policy improvement in our experimental setups. The planning horizon $T$ is determined dynamically from the distribution of $\Phi^{k}_{t+l}$. For each prediction step $l$, we check whether the distribution of $\Phi^{k}_{t+l}$ is close to uniform (according to Kullback-Leibler distance) or whether any of states with positive reward has probability above a threshold; in that case, the prediction cycle is stopped. The action value is determined by combining the corresponding SF and feature state rewards as shown in equation 11. Finally, the action is sampled from the softmax distribution over the action values.
@@ 3.5 DHTM and Episodic Control
DTHM performs best when its hidden state space is large. Since the hidden state for unpredicted observations is formed randomly, it’s important that there are no accidental state collisions when remembering transitions. That is, DHTM tries to remember new sequences without interfering with old memories. In effect, this makes DHTM a buffer of agent trajectories with a built-in recall engine. Episodic Control is based on a similar idea, and to investigate its difference from DHTM, we implemented an Episodic Control agent, a simplified counterpart of DHTM, reflecting its core property that presumably helps it solve RL tasks.
Episodic Control (EC) is a general computational approach to storing and reusing an agent’s experience, inspired by the notion of episodic memory from psychology (Tulving et al., 1972). The core of EC is a dictionary $\mathcal{D}_{a}=(K_{a},V_{a})$ that stores key-value pairs for each discrete action $a\in A$. In our EC version, $K_{a}$ is an array of hidden states $h_{t}$ and $V_{a}$ is an array of corresponding subsequent hidden states $h_{t+1}$. If the key state $h_{t}$ already exists in the dictionary, its value is rewritten. That is, $\mathcal{D}_{a}$ represents the transition matrix of a deterministic HMM for a feature variable $\Phi_{t}$. The hidden state $h_{t}$ and the feature state $\varphi_{t}$ have the same dependence as in DHTM (equation 6), but in EC we assume that there are infinite hidden states per feature state. To predict the next state, we just have to look up $\mathcal{D}_{a}$ for the current state $h_{t}$, but if there is no entry for $h_{t}$ or the prediction does not match the observed state $\varphi_{t+1}$, then the new state is formed. Unlike DHTM, the new state is not completely random, but to avoid collisions, the algorithm makes sure that this state hasn’t been chosen before. The learning procedure of the EC agent’s memory $\mathcal{D}_{a}$ is summarized in the algorithm 2.
Since $\mathcal{D}_{a}$ can be transformed into a transition matrix, the process of forming SF can be reduced to one in Sec. 3.4. However, since the resulting transition matrix is sparse, it’s more efficient to consider $\mathcal{D}_{a}$ as a graph representation and to perform SF formation as a breadth-first search (see Appendix E.3 for details).
@@ 4 Experiments
We test our model in a reinforcement learning task in the Gridworld environment and in a more challenging AnimalAI 3D environment. This section shows how different memory models affect the adaptability of an RL agent. For each setup, the results are averaged over five experiments with different seed values, and one standard deviation is shown everywhere, except for boxplots. See Appendix E for details on the baselines, their model parameters, and training regimes. See Appendix F for a more detailed description of the environments and setups [^1] The source code of the experiments and algorithms is at https://github.com/Cognitive-AI-Systems/him-agent/tree/iclr25..
@@ 4.1 Gridworld
Figure 2: POMDP 5x5 Gridworld foraging task using different types of memory: DHTM, CSCG, LSTM, RWKV. All agents, except for DHTM, are trained in two modes. The first mode is with oracle, which tells the agent to reinitialize memory and clear the trajectory buffer during environmental change. The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories and update memory. There is also an episodic control agent (ec) among baselines, which uses a simple dictionary to store trajectories and make predictions. A. Dynamic of steps required for the agent to reach the goal. In the 300th episode, the goal is blocked by a wall, so the agent needs to find a new optimal trajectory. B. Total steps taken during 1000 episodes.
@@
The first test was conducted in partially observable Gridworld in a 5x5 alternating maze with a stationary end state and the agent’s starting position. The agent can observe only the colour of the cell it is in and the colour of an obstacle it encounters, while being punished in addition to the constant step penalty. Each episode ends when the reward is collected or when the maximum number of action steps is reached. In each episode, we measure the number of action steps required for the agent to reach the goal.
@@
Each trial lasts $1000$ episodes and consists of two phases. In the first phase, there are no obstacles on the shortest path to the terminal state. In the second phase, after $300$ episodes, the previously optimal path is blocked by a wall, so the agent has to take a detour to get the reward. The experiment aims to show the ability of a memory to relearn the structure of the maze in the test time. If the agent is able to reach the end state in the second phase, it suggests that the model can forget, or at least suppress, the memory of the old shortest path when it is blocked. Otherwise, the agent will continue to hit the wall and fail to reach the terminal state.
@@
In this experiment, we compare our memory model DHTM with CSCG (George et al., 2021), LSTM (Hochreiter & Schmidhuber, 1997) and RWKV (Peng et al., 2023) within the same agent architecture that uses memory to form successor features and evaluate actions (see Sec. 3.4). All memories, except DHTM, are trained in two modes. The first mode is with an oracle that tells the agent to reinitialize memory and clear the trajectory buffer just in time for the environmental change. The second mode is without an oracle, so the agent continues to use its current strategy to collect new trajectories and update memory in the second phase. There is also a simple episodic control agent (EC) among the baselines, which uses a dictionary to store trajectories and make predictions (see Section 3.5 for details).
@@
The results in Figure 2 show that CSCG effectively can’t adapt to the environmental change in the online regime, but only if it is retrained from scratch on new trajectories. And, apparently, LSTM and RWKV are only able to learn a predictive model for a small horizon, since they perform significantly worse than other baselines in the second phase, even with oracle enabled. In contrast, DHTM and EC are able to adapt online even without explicit memory’s trainable parameters reinitialisation. Another important observation is that the EC agent performs on par with DHTM, suggesting that its mechanism can explain most of DHTM’s performance in this task. However, see Appendix H.2 for experiments that show their difference in scalability.
@@ 4.2 AnimalAI
Figure 3: Amount of action steps the agent performs to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM, CSCG, LSTM, RWKV and EC dictionary. The agent and the goal have the same initial positions throughout the episodes. Each time-step, the agent gets an encoded image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward. A. Dynamic of steps required for the agent to reach the goal. B. Total steps taken during 1000 episodes by the agent with DHTM with K-Means encoder and Categorical VAE. For the latter, DHTM’s context factor size is varied (i.e., how many previous time-step hidden variables are considered for prediction).
@@
We test DHTM and EC agents using two different encoders: K-Means and Categorical VAE (Jang et al., 2017). Both encoders are pre-trained on RGB images collected with a uniform strategy during $2000$ episodes in the same room used for agent testing. K-Means represents an image as a single categorical variable with $50$ states, and Categorical VAE is configured for distributed representation with five variables with $50$ states each (see Appendix D for details). In case of EC agent, since it doesn’t support distributed representations, we map each unique VAE encoding to one feature state using a simple dictionary. Since the resulting number of states for such feature variable is large (typically more than $1000$), it’s infeasible to use the same mapping for CSCG. We also didn’t find much difference for LSTM and RWKV using K-Means or VAE, so we only keep the results for VAE encoder.
@@
To show that DHTM can’t simply be replaced by multiple parallel EC dictionaries or CSCGs in the case of distributed encoding, we also vary the context factor size of DHTM, that is, how many previous time-step variables are taken into account to predict the current state of a variable. The factor size of 1 corresponds to the graph of the Factorial HMM, with parallel HMMs. It can be seen from Figure 3B that in our case of categorical VAE, knowing only the previous state of the variable is not enough to reliably predict rewarding features, since the agent doesn’t converge to the shortest path. We also tested the variant (factor 1*) where each feature variable is represented by three hidden states, to ensure that the result for factor size $1$ is not due to DHTM’s hidden state collisions. DHTM with a factor that takes into account more than one previous time-step hidden variable (excluding the action variable) performs significantly better. For diagrams depicting factor graphs used in these experiments, see Appendix E.4. For further experiments showing the importance of distributed encoding in DHTM, see Appendix H.3.
@@ 5 Conclusion
Our results suggest that DHTM can be viewed as an adaptive trajectory buffer for a slower generalizing algorithm, such as a Neural State Space Model (Jiang et al., 2023). Through prediction error learning, DHTM can selectively discard and replace obsolete transitions in observed trajectories without disrupting relevant experience. It can also provide an initial policy for an agent through episodic control until a slower but more accurate world model takes over. These features of a trajectory buffer are highly desirable in the case of online learning in non-stationary environments.
@@ Appendix A Value Function Decomposition
Appendix B Details of Neuronal Implementation
As shown in Figure 1B, each RV can be viewed as a set of neurons representing the states of the RV, i.e. $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$. Cell activity is binary $c^{j}_{t}\in\{0,1\}$, and the probability can be interpreted as a spike rate. Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor. Let’s denote the elements of the vectors as $f_{l}$ and $m_{u}$ respectively, where $l$ corresponds to a particular combination of state values $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$, and $u$ indexes all neurons representing states of RVs from previous time steps.
We can now rewrite equation 3 as the following:
$p(h_{t}^{k})\propto\sum_{l\in\mathrm{seg}(j)}L_{l}f^{k}_{l},$ (14)
where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column. The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM (George et al., 2021).
where $w_{ul}$—synapse efficacy or neuron specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field $rec(l)$.
where ${E_{l}=\log f_{l}+\log L_{l}}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells representing states for the variable $H^{k}_{t}$. Here we also approximate the logarithmic sum with the $\mathrm{max}$ operation, inspired by the neurophysiological model of segment aggregation by cells (Stuart & Spruston, 2015).
Appendix C Biological Interpretation
Neural implementation of the DHTM is inspired by neocortical neural networks (see Fig. 4). Hidden variables $H^{k}$ may be considered as populations of excitatory pyramidal neurons in cortical layer L2/3 of somatosensory areas, with lateral inhibition modelled as $\mathrm{softmax}$ function. Staiger & Petersen (2021) showed that neurons in this layer are responsible for temporal context formation.
The neuronal activity at timestep $t$ can be thought to carry messages $M^{k}_{t-1}$. Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$. Dendritic segments of biological neurons are known to be coincidence detectors of their synaptic input (Stuart & Spruston, 2015). That is, each segment can be thought of as a detector of a particular combination of active cells (states in our interpretation). If we associate a segment’s parameter with factor value, a collection of dendritic segments represent a mapping from state space to factor values, representing context factor $F_{c}$.
Our computational model of neuron and its dendrites is based on HTM neuron model (Hawkins & Ahmad, 2016). In this model, each segment has a receptive field and activates when a number of active cells in this receptive field reaches a threshold. The neuron is active if any of its segments are active. We use the same activation rule, but the threshold is fixed to the receptive field size. Our algorithm of segment’s growth on demand is also similar to HTM’s. In principle, DHTM can be considired as a probabilistic interpretation of HTM, allowing belief propagation.
Similarly to HTM framework, our temporal memory can be interpreted as a model of cortical layers. Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3. L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ (Mountcastle, 1997).